OpenStudy (anonymous):

Which proportion models this problem? On a soccer team, 5 out of 13 players surveyed said they had two or more siblings. The league has 520 players. Which is the best prediction of the number of players in the league that have two or more siblings?

OpenStudy (p0sitr0n):

5/13*520=2600/13

OpenStudy (p0sitr0n):

2600/13=200 players